# Transition from topological to chaos in the nonlinear Su-Schrieffer-Heeger model

### v2

@@ Abstract
Next, we consider the even numbers of rows. Each of them becomes $\tilde{a}(\Psi_{A}(x),\Psi_{B}(x))\Psi_{A}(x)=d\Psi_{A}(x+1)$. Therefore, we obtain $\Psi_{A}(x+1)=\tilde{a}(\Psi_{A}(x),\Psi_{B}(x))\Psi_{A}(x)/d=[a+b(|\Psi_{A}(x)|^{2}+|\Psi_{B}(x)|^{2}]\Psi_{A}(x)/d$, which is equivalent to Eq. (19).
.3 Numerical calculation of the bifurcation diagram of the nonlinear SSH model.
@@
To obtain the bifurcation diagram in Fig. 2, we numerically calculate the nonlinear discrete dynamics in Eq. (19). We here fix the parameters $b=-1$ and $d=1$ and change the parameter $a$. At any $a$, we set the initial condition as $\Psi_{A}(1)=0.1$. To obtain $\Psi_{A}(x)$ at $x>1$, we iteratively calculate the right-hand side of Eq. (19) and update the value of $\Psi_{A}(x+1)$ by using $\Psi_{A}(x)$. We calculate $\Psi_{A}(x)$ for $1\leq x\leq 10000$ at each $a$. Then, we plot the values of $\Psi_{A}(x)$ of $9900<x\leq 10000$ in Fig. 2.
@@
.4 Lyapunov exponents.
Lyapunov exponents are the rate of the amplification or attenuation of the perturbation from the solution of a nonlinear dynamical system, whose positivity indicates the chaos of the dynamical system. We numerically confirm the bifurcation to chaos by calculating the Lyapunov exponent of the dynamical system describing zero modes in the nonlinear SSH model (Eq. (7) in the main text). One can calculate the Lyapunov exponent of a one-dimensional discrete dynamical system $\psi_{i+1}=f(\psi_{i})$ such as Eq. (7) in the main text from the following formula:
@@
Supplementary Figure S1 shows the numerically obtained Lyapunov exponents of the dynamical system describing zero modes in the nonlinear SSH model. We use $L=10000$ and $b=c=d=1$. Around $a=2.3$, the Lyapunov exponent becomes positive, which indicates the bifurcation to chaos. As is usually seen in chaotic maps, even larger $a$ than the chaos transition point, Lyapunov exponents can be negative in some parameter regions. However, in these parameter regions, the steady-state solutions are periodic and thus still break the bulk-edge correspondence.
.5 Bulk-edge correspondence in finite systems.
We here numerically confirm the bulk-edge correspondence in finite systems of the nonlinear SSH model (Eqs. (4) and (5) in the main text). To judge the localization or anti-localization of zero modes, we define the indicator of localization $P$ as
@@
with $L$ being the system size. $\Psi_{j}(x)$ is the nonlinear eigenvector of a zero mode and $x$ and $j$ represent the location and the internal degrees of freedom, respectively. This localization indicator becomes larger (smaller) than one, when the zero mode is localized (anti-localized) at the edge. We note that one should modify the definition of the localization indicator in long-range models, while we here focus on the nonlinear SSH model whose hopping range is one.
@@
Figure S2: Minimum absolute values of eigenvalues and localization indicators $P$ of the finite system of the nonlinear SSH model. We fix the parameters $b=c=d=1$. We plot the minimum absolute values of the nonlinear eigenvalues at the system size $L=5$, $10$, $15$, and $20$ from the top. The red (black) circles represent the localization $P>1$ (delocalization $P<1$) of the corresponding nonlinear eigenvectors. a Localization indicators at $a=1.5$. The blue dashed lines are the transition point of the nonlinear winding number. We can confirm the bulk-edge correspondence between the nonlinear winding number and the localized zero modes. We note that some data points disappear at $a>2.5$ due to the difficulty of the convergence of the numerical techniques under the strong nonlinearity. b Localization indicators at $a=1.9$. The localization and delocalization of zero modes are switched at the transition point represented by the blue dashed line, which indicates the bulk-edge correspondence. c Localization indicators at $a=2.1$. Even at the amplitude larger than the critical amplitude $w=1.1$ (the blue dashed line), there exist anti-localized zero modes. Therefore, the bulk-edge correspondence is broken in this parameter regime, which is consistent with the analysis in the semi-infinite system.
@@
Figure S3: Size- and nonlinearity-dependence of the minimum absolute values of nonlinear eigenvalues. a Finite-size scaling of the minimum absolute values of nonlinear eigenvalues in the nonlinear SSH model. We use the parameters $a=0.6$ and $b=c=d=1$. We confirm that the minimum eigenvalues are inversely proportional to the system size and thus converge to zero in the thermodynamic limit. b Disappearance of the anti-localized zero mode in the linear limit. The red dashed line shows the bulk band gap obtained from the Bloch ansatz. The horizontal axis shows the value of $b$ (we set $b=c$). We fix the other parameters as $a=1.5$, $d=1$, and the amplitude $w=0.1$. At $b=0.3$, the minimum absolute values of the nonlinear eigenvalues become larger than the bulk band gap. Thus, one cannot distinguish the anti-localized modes and bulk modes at $b<0.3$, which implies the disappearance of the anti-localized zero modes in the linear limit.
@@
We numerically solve the nonlinear eigenvalue problem of the finite chains of the nonlinear SSH model with different system sizes $L=5$, $10$, $15$, and $20$. To solve the nonlinear eigenvalue problem, we assume the eigenequation as an algebraic equation and use the quasi-Newton method Sone2023 . Supplementary Figure S2 shows the minimum absolute values of eigenvalues and the localization indicators (Supplementary Eq. (S59)) of the corresponding eigenvectors. At $a=1.5$ (Supplementary Fig. S2a), the nonlinear winding number is nonzero $\nu_{\rm NL}=1$ in the range of amplitude, $2.50.5<w<$ 2.5. In this regime, we obtain localized zero modes corresponding to the nonzero winding number. Meanwhile, we obtain anti-localized zero modes at $w<0.5$ and no zero modes at $w>2.5$. Therefore, the nonlinear winding number predicts the existence or absence of the localized zero modes at $a=1.5$. We can also confirm the bulk-edge correspondence at $a=1.9$ as shown in Supplementary Fig. S2b.
@@
In contrast, if we consider $a=2.1$, we obtain anti-localized zero modes even in the topologically nontrivial phase at $w>1.1$ (Supplementary Fig. S2c). Therefore, the bulk-edge correspondence is broken in this parameter regime, which is consistent with the result in the semi-infinite system shown in Fig. 2 in the main text. We note that the zero modes in the semi-infinite system exhibit the localization indicator $P>1$ even at $a>2$ and thus can be regarded as locally localized but globally delocalized zero modes. Meanwhile, the zero modes in finite systems can be locally delocalized in the sense that the localization indicator $P$ is smaller than one. While both the breakdowns of the bulk-edge correspondence observed in finite and semi-infinite systems are related to the period-doubling bifurcation, the detailed mechanism of the delocalization of zero modes in finite systems remains a future issue.
@@
The obtained nonlinear eigenvalues are not exactly zero due to the finite-size effect. To confirm the convergence of nonlinear eigenvalues to zero in the thermodynamic limit $L\rightarrow\infty$, we investigate the finite-size scaling of the minimum absolute values of nonlinear eigenvalues. Supplementary Figure S3a shows the minimum absolute values of nonlinear eigenvalues at different system sizes. We here use the parameters $a=1.5$, $b=c=1$, and $d=1$, and fix the edge amplitude to be $w_{\rm edge}=0.6$. We confirm that the nonlinear eigenvalues are inversely proportional to the system size and thus converge to zero in the thermodynamic limit.
@@
It is also noteworthy that the anti-localized zero modes disappear in the linear limit $b,c\rightarrow 0$. We consider the case of $b=c$ and numerically confirm such disappearance of the anti-localized zero modes by calculating the nonlinear eigenvalues at different $b$. Supplementary Figure S3b shows the minimum absolute values of the nonlinear eigenvalues and the size of the bulk band gaps predicted from the Bloch ansatz. At sufficiently small $b$, the nonlinear eigenvalues become larger than the bulk band gaps, and thus one cannot distinguish the anti-localized zero modes and bulk modes. Therefore, anti-localized zero modes are induced by the nonlinearity and disappear in the linear limit.
@@
Figure S4: Minimum absolute values of eigenvalues and different definitions of localization indicators $P_{1,2,3}$ of the finite system of the nonlinear SSH model. We fix the parameters $b=c=d=1$. We plot the minimum absolute values of the nonlinear eigenvalues at the system size $L=5$, $10$, $15$, and $20$ from the top. The red (resp. black) circles represent $P_{1,2,3}>1$ (resp. $P_{1,2,3}<1$) of the corresponding nonlinear eigenvectors. a-c Localization indicators at $a=1.5$. Panels a, b, and c show the localization indicators $P_{1}$, $P_{2}$, and $P_{3}$, respectively. The blue dashed lines are the transition point of the nonlinear winding number. We can confirm the bulk-edge correspondence between the nonlinear winding number and the localized zero modes. We note that some data points disappear at an amplitude larger than $2.5$ due to the difficulty of the convergence of the numerical techniques under the strong nonlinearity. In this parameter region, all the indicators correspond to the bulk topological invariants. d-f Localization indicators at $a=1.9$. Panels d, e, and f show the localization indicators $P_{1}$, $P_{2}$, and $P_{3}$, respectively. The blue dashed lines show the transition point of the nonlinear winding number, while the localization indicators are unchanged at those boundaries. Since the bulk-edge correspondence should still be valid in this parameter region as inferred from the bifurcation diagram in Fig. 2 in the main text, the localization indicators $P_{1,2,3}$ are useless to discuss the bulk-boundary boundary correspondence in finite systems.

### v3

@@ Abstract
Next, we consider the even numbers of rows. Each of them becomes $\tilde{a}(\Psi_{A}(x),\Psi_{B}(x))\Psi_{A}(x)=d\Psi_{A}(x+1)$. Therefore, we obtain $\Psi_{A}(x+1)=\tilde{a}(\Psi_{A}(x),\Psi_{B}(x))\Psi_{A}(x)/d=[a+b(|\Psi_{A}(x)|^{2}+|\Psi_{B}(x)|^{2}]\Psi_{A}(x)/d$, which is equivalent to Eq. (19). Based on the real–space description of the nonlinear eigenequation (Eq. (S57)) and its reduction (Eq. (19)), we calculate zero modes of the nonlinear SSH model. We note that this spatial dynamics is different from the temporal dynamics in Eqs. (Transition from topological to chaos in the nonlinear Su-Schrieffer-Heeger model) and (11), and we do not calculate zero modes from the steady state of the temporal dynamics.
.3 Cubic map and its correspondence to the spatial dynamics in Eq. (19).
The cubic map is a nonlinear discrete dynamics defined as
$\Phi(t+1)=(1-a^{\prime})\Phi(t)+a^{\prime}\Phi(t)^{3},$ (S58)
where $\Psi(t)$ is the state variable at time $t$. This discrete dynamics exhibits a chaos transition at $a^{\prime}\sim 3.3$ Rogers1983 . One can confirm such a chaos transition from the positivity of the Lyapunov exponent. The nonlinear dynamics in Eq. (19) is equivalent to the cubic map by transforming $a$ and $\Psi(x)$ as $a=(a^{\prime}-1)d$ and $\Psi(x)=(-1)^{x}\sqrt{a^{\prime}d/b}\Phi(x)$ and assuming the space index $x$ as time $t$. The chaos transition point in Fig. 2 is consistent with the previous study on the cubic map.
.4 Numerical calculation of the bifurcation diagram of the nonlinear SSH model.
@@
To obtain the bifurcation diagram in Fig. 2, we numerically calculate the nonlinear discrete dynamics in Eq. (19). We here fix the parameters $b=-1$ and $d=1$ and change the parameter $a$. At any $a$, we set the initial condition as $\Psi_{A}(1)=0.1$. To obtain $\Psi_{A}(x)$ at $x>1$, we iteratively calculate the right-hand side of Eq. (19) and update the value of $\Psi_{A}(x+1)$ by using $\Psi_{A}(x)$. We calculate $\Psi_{A}(x)$ for $1\leq x\leq 10000$ at each $a$. Then, we plot the values of $\Psi_{A}(x)$ of $9900<x\leq 10000$ in Fig. 2. We have also used the same methods in Fig. 3.
.5 Numerical calculation of the vector field.
We plot the vector field $(\Psi_{A}(x+1)-\Psi_{A}(x),\Psi_{A}(x+2)-\Psi_{A}(x+1))$ in Fig. 4 by using the StreamPlot function in Mathematica.
.6 Reduction of a higher–dimensional model into a one–dimensional model.
As is discussed in the main text, higher–dimensional models of nonlinear topological insulators are often reduced to one–dimensional models, when we consider specific wavenumbers, such as $k_{y}=0,\pi$. Therefore, the technique of dynamical systems used in our analysis can be extended to any dimensions, for which our result should provide a guiding principle to investigate the bulk–edge correspondence in nonlinear systems.
We here explicitly show that a two–dimensional model termed the nonlinear Qi–Wu–Zhang (QWZ) model Sone2023 is reduced to the nonlinear SSH model in this article. The nonlinear eigenvalue problem of the nonlinear QWZ model is described in the wavenumber space as
$\left(\begin{array}[]{cc}u+\kappa w(\psi)+\cos k_{x}+\cos k_{y}&\sin k_{x}+i\sin k_{y}\\ \sin k_{x}-i\sin k_{y}&-[u+\kappa w(\psi)+\cos k_{x}+\cos k_{y}]\end{array}\right)\left(\begin{array}[]{c}\psi_{A}(\mathbf{k})\\ \psi_{B}(\mathbf{k})\end{array}\right)=E\left(\begin{array}[]{c}\psi_{A}(\mathbf{k})\\ \psi_{B}(\mathbf{k})\end{array}\right),$ (S65)
where $(\psi_{A}(\mathbf{k}),\psi_{B}(\mathbf{k}))^{T}$ is the nonlinear eigenvector at the wavenumber $\mathbf{k}=(k_{x},k_{y})$, and $w(\psi)=|\psi_{A}(\mathbf{k})|^{2}+|\psi_{B}(\mathbf{k})|^{2}$ represents the nonlinear term in the nonlinear QWZ model. When we focus on the wavenumber $k_{y}=0$, the nonlinear eigenvalue problem reads
$\left(\begin{array}[]{cc}u+1+\kappa w(\psi)+\cos k_{x}&\sin k_{x}\\ \sin k_{x}&-[u+1+\kappa w(\psi)+\cos k_{x}]\end{array}\right)\left(\begin{array}[]{c}\psi_{A}(k_{x},0)\\ \psi_{B}(k_{x},0)\end{array}\right)=E\left(\begin{array}[]{c}\psi_{A}(k_{x},0)\\ \psi_{B}(k_{x},0)\end{array}\right).$ (S72)
Furthermore, if we consider a unitary transformation of the nonlinear eigenvector
$\left(\begin{array}[]{c}\tilde{\psi}_{A}(k_{x},0)\\ \tilde{\psi}_{B}(k_{x},0)\end{array}\right)=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}\psi_{A}(k_{x},0)-i\psi_{B}(k_{x},0)\\ \psi_{A}(k_{x},0)+i\psi_{B}(k_{x},0)\end{array}\right),$ (S73)
Eq. (S72) becomes
$\left(\begin{array}[]{cc}0&u+1+\kappa w(\tilde{\psi})+e^{-ik_{x}}\\ u+1+\kappa w(\tilde{\psi})+e^{ik_{x}}&0\end{array}\right)\left(\begin{array}[]{c}\tilde{\psi}_{A}(k_{x},0)\\ \tilde{\psi}_{B}(k_{x},0)\end{array}\right)=E\left(\begin{array}[]{c}\tilde{\psi}_{A}(k_{x},0)\\ \tilde{\psi}_{B}(k_{x},0)\end{array}\right),$ (S80)
with $w(\tilde{\psi})$ being $w(\tilde{\psi})=|\tilde{\psi}_{A}(k_{x},0)|^{2}+|\tilde{\psi}_{B}(k_{x},0)|^{2}$. Rewriting the parameters $u$, $\kappa$ as $a=u+1$, $b=\kappa$, the above eigenequation is equivalent to that of the nonlinear SSH model (Eq. (18)). A similar reduction is conducted at $k_{y}=\pi$.
For general $k_{y}$, the unitary transformation in Eq. (S73) modifies the nonlinear eigenvalue problem in Eq. (S65) as
$\left(\begin{array}[]{cc}\sin k_{y}&u^{\prime}+\kappa w(\tilde{\psi})+e^{-ik_{x}}\\ u^{\prime}+\kappa w(\tilde{\psi})+e^{ik_{x}}&-\sin k_{y}\end{array}\right)\left(\begin{array}[]{c}\tilde{\psi}_{A}(k_{x},k_{y})\\ \tilde{\psi}_{B}(k_{x},k_{y})\end{array}\right)=E\left(\begin{array}[]{c}\tilde{\psi}_{A}(k_{x},k_{y})\\ \tilde{\psi}_{B}(k_{x},k_{y})\end{array}\right),$ (S87)
with $u^{\prime}$ being $u^{\prime}=u+\cos k_{y}$. Then, by fixing $k_{y}$, we obtain an effectively one–dimensional nonlinear eigenequation in the real space as
$E\Psi_{A}(x)$ $=$ $\sin k_{y}\Psi_{A}(x)+(u^{\prime}+\kappa|\Psi_{A}(x)|^{2}+\kappa|\Psi_{B}(x)|^{2})\Psi_{B}(x)+\Psi_{B}(x-1),$ (S88)
$E\Psi_{B}(x)$ $=$ $-\sin k_{y}\Psi_{B}(x)+(u^{\prime}+\kappa|\Psi_{A}(x)|^{2}+\kappa|\Psi_{B}(x)|^{2})\Psi_{A}(x)+\Psi_{A}(x+1),$ (S89)
which is a nonlinear SSH model with staggered on–site potential. The eigenvalues of nonlinear edge modes in this effectively one–dimensional system are estimated as $E=\sin k_{y}$ for left–localized modes and $E=-\sin k_{y}$ for right–localized modes. Assuming such dispersion relation of the edge mode, we obtain a nonlinear dynamical system similar to Eq. (19). Therefore, in the nonlinear QWZ model, the chaos transition of edge modes can occur at arbitrary $k_{y}$. In general models, specifying the dispersion relation of edge modes can be a nontrivial problem, and thus we have to develop a way to determine the eigenvalue and eigenvector at the same time.
.7 Possible optical setup of the nonlinear SSH model.
The nonlinear SSH model (Eqs. (4) and (5) in the main text) can describe various systems, such as topological photonics and electrical circuits. We here discuss the possible realization of the nonlinear coupling used in the model by using photonic systems. Specifically, we consider an interaction between two sites via an optical fiber used in Ref. Bisianov2019 . When one uses such an optical fiber, the phase of the coupling depends on the amplitude of the light because the light propagating in the optical fiber is affected by the Kerr nonlinearity, and thus the phase shift occurs depending on its amplitude. Thus, the coupling term has a form of $ce^{i\kappa|\Psi_{j}|^{2}}\Psi_{j}$ with $c$ being the strength of the coupling and $\kappa$ being the coefficient of the Kerr nonlinearity. Remaining the leading order terms, we obtain $c(1+i\kappa|\Psi_{j}|^{2}-\kappa^{2}|\Psi_{j}|^{4}/2+\cdots)\Psi_{j}$. The second term has a similar form to the nonlinear coupling in the nonlinear SSH model except for the multiplication of the imaginary unit $i$. The third term is a higher–order term without the imaginary unit and thus can induce the nonlinearity–induced topological phase transition as in the models discussed in the main text.
@@
.8 Lyapunov exponents.
Lyapunov exponents are the rate of the amplification or attenuation of the perturbation from the solution of a nonlinear dynamical system, whose positivity indicates the chaos of the dynamical system. We numerically confirm the bifurcation to chaos by calculating the Lyapunov exponent of the dynamical system describing zero modes in the nonlinear SSH model (Eq. (7) in the main text). One can calculate the Lyapunov exponent of a one–dimensional discrete dynamical system $\psi_{i+1}=F(\psi_{i})$ such as Eq. (7) in the main text from the following formula:
@@
Supplementary Figure S1 shows the numerically obtained Lyapunov exponents of the dynamical system describing zero modes in the nonlinear SSH model. We use $L=10000$ and $b=c=d=1$. Around $a=2.3$, the Lyapunov exponent becomes positive, which indicates the bifurcation to chaos. We estimate the chaos transition point in Fig. 2 in the main text from this numerical result. As is usually seen in chaotic maps, even larger $a$ than the chaos transition point, Lyapunov exponents can be negative in some parameter regions. However, in these parameter regions, the steady–state solutions are periodic and thus still break the bulk–edge correspondence.
One can also find zero Lyapunov exponents at some $a$’s. Such $a$’s correspond to the period–doubling bifurcation points. In particular, $a=1$ and $a=2$ are the bifurcation points as presented in Fig. 2 in the main text. Zero Lyapunov exponents at $a=1$ and $a=2$ can be analytically confirmed from the fact that the corresponding fixed points are linearly marginal in the spatial dynamics (Eq. (7) in the main text), i.e., the derivative of $F$ is zero at the fixed points: $|F^{\prime}(a=1,\tilde{\psi}_{i}=0)|=0$ and $|F^{\prime}(a=2,\tilde{\psi}_{i}=1)|=0$.
.9 Temporal linear stability and instability of zero modes.
While we have analyzed the stability of the spatial dynamics of zero modes in the previous section, we also conduct the linear stability analysis of the temporal dynamics of zero modes. To this end, we consider the linearization of Eqs. (4) and (5) in the main text around the periodic solution $\Psi^{0}_{A,B}(x)=\psi_{A,B}e^{ikx}$. We obtain the linearized dynamics,
$i\frac{\partial}{\partial t}\left(\begin{array}[]{c}\delta\psi_{A}\\ \delta\psi_{B}\\ \delta\psi_{A}^{\ast}\\ \delta\psi_{B}^{\ast}\end{array}\right)$ $=$ $H_{L}(\psi_{A},\psi_{B})\left(\begin{array}[]{c}\delta\psi_{A}\\ \delta\psi_{B}\\ \delta\psi_{A}^{\ast}\\ \delta\psi_{B}^{\ast}\end{array}\right)$ (S99)
$=$ $\left(\begin{array}[]{cccc}b\psi_{A}^{\ast}\psi_{B}&\tilde{a}b|\psi_{B}|^{2}&b\psi_{A}\psi_{B}&b\psi_{B}^{2}\\ \tilde{a}^{\ast}+b|\psi_{A}|^{2}&b\psi_{A}\psi_{B}^{\ast}&b\psi_{A}^{2}&b\psi_{A}\psi_{B}\\ -b\psi_{A}^{\ast}\psi_{B}^{\ast}&-b\psi_{B}^{\ast 2}&-b\psi_{A}\psi_{B}^{\ast}&-\tilde{a}^{\ast}-b|\psi_{B}|^{2}\\ -b\psi_{A}^{\ast 2}&-b\psi_{A}^{\ast}\psi_{B}^{\ast}&-\tilde{a}-b|\psi_{A}|^{2}&-b\psi_{A}^{\ast}\psi_{B}\end{array}\right)\left(\begin{array}[]{c}\delta\psi_{A}\\ \delta\psi_{B}\\ \delta\psi_{A}^{\ast}\\ \delta\psi_{B}^{\ast}\end{array}\right),$ (S108)
where we consider a periodic modulation $\Psi_{A,B}(x)=\Psi^{0}_{A,B}(x)+\delta\psi_{A,B}e^{ik^{\prime}x}$ and $\tilde{a}=a-b(|\psi_{A}|^{2}+|\psi_{B}|^{2})+de^{ik^{\prime}}$. We note that the wavenumber of this periodic modulation can be different from that of the periodic solution $\Psi^{0}_{A,B}(x)$.
By calculating the eigenvalues of $H_{L}(\psi_{A},\psi_{B})$, we find that the bifurcation in the nonlinear dynamical system in the spatial direction (Eq. (7) in the main text) is also related to the temporal instability of zero modes. We here consider the periodic solution corresponding to bulk modes with zero eigenvalues: $k=\pi$, $(\psi_{A},\psi_{B})=(w/\sqrt{2},\pm w/\sqrt{2})$, $w=a-1$. Then, the eigenvalues of $H_{L}(\psi_{A},\psi_{B})$ become $\lambda_{1\pm}=\pm\sqrt{|a+de^{ik^{\prime}}|^{2}+2bw(a+d\cos k^{\prime})+(bw)^{2}}$ and $\lambda_{2\pm}=\pm\sqrt{|a+de^{ik^{\prime}}|^{2}+4bw(a+d\cos k^{\prime})+3(bw)^{2}}$. $\lambda_{1\pm}$ are always real, while $\lambda_{2\pm}$ become purely imaginary if $|a+de^{ik^{\prime}}|^{2}+4bw(a+d\cos k^{\prime})+3(bw)^{2}$ is negative. Since $|a+de^{ik^{\prime}}|^{2}+4bw(a+d\cos k^{\prime})+3(bw)^{2}$ takes minimum at $k^{\prime}=0$ or $k^{\prime}=\pi$ and becomes zero at $k^{\prime}=\pi$, $\lambda_{2\pm}$ are real if $(a+d)^{2}+4bw(a+d)+3(bw)^{2}$ is positive. In the case of negative $b$, we obtain the condition for the linear instability, i.e., imaginary $\lambda_{2\pm}$ as $(a+d)/(3|b|)<w<(a+d)/|b|$. When we assume the parameters $a>0$, $b=-1$, and $d=1$ as in Fig. 2 in the main text, one can confirm that this inequality is equivalent to $a>2$, at which the localized zero mode becomes unstable in the spatial dynamics in Eq. (7) in the main text. Thus, the spatial and temporal instability of zero modes are closely related to each other.
.10 Bulk-edge correspondence in finite systems.
We here numerically confirm the bulk–edge correspondence in finite systems of the nonlinear SSH model (Eqs. (4) and (5) in the main text). To judge the localization or anti–localization of zero modes, we define the indicator of localization $P$ as
@@
with $L$ being the system size. $\Psi_{j}(x)$ is the nonlinear eigenvector of a zero mode and $x$ and $j$ represent the location and the internal degrees of freedom, respectively. This localization indicator becomes larger (smaller) than one, when the zero mode is localized (anti–localized) at the edge. We note that one should modify the definition of the localization indicator in long–range models, while we here focus on the nonlinear SSH model whose hopping range is one.
@@
Figure S2: Minimum absolute values of eigenvalues and localization indicators $P$ of the finite system of the nonlinear SSH model. We fix the parameters $b=c=d=1$. We plot the minimum absolute values of the nonlinear eigenvalues at the system size $L=5$, $10$, $15$, and $20$ from the top. The red (black) circles represent the localization $P>1$ (anti–localization $P<1$) of the corresponding nonlinear eigenvectors. a Localization indicators at $a=1.5$. The blue dashed lines are the transition point of the nonlinear winding number. We can confirm the bulk–edge correspondence between the nonlinear winding number and the localized zero modes. We note that some data points disappear at $a>2.5$ due to the difficulty of the convergence of the numerical techniques under the strong nonlinearity. b Localization indicators at $a=1.9$. The localization and anti–localization of zero modes are switched at the transition point represented by the blue dashed line, which indicates the bulk–edge correspondence. c Localization indicators at $a=2.1$. Even at the amplitude larger than the critical amplitude $w=1.1$ (the blue dashed line), there exist anti–localized zero modes. Therefore, the bulk–edge correspondence is broken in this parameter regime, which is consistent with the analysis in the semi–infinite system.
@@
Figure S3: Size– and nonlinearity–dependence of the minimum absolute values of nonlinear eigenvalues. a Finite-size scaling of the minimum absolute values of nonlinear eigenvalues in the nonlinear SSH model. We use the parameters $a=0.6$ and $b=c=d=1$. We confirm that the minimum eigenvalues are inversely proportional to the system size and thus converge to zero in the thermodynamic limit. b Disappearance of the anti–localized zero mode in the linear limit. The red dashed line shows the bulk band gap obtained from the Bloch ansatz. The horizontal axis shows the value of $b$ (we set $b=c$). We fix the other parameters as $a=1.5$, $d=1$, and the amplitude $w=0.1$. At $b=0.3$, the minimum absolute values of the nonlinear eigenvalues become larger than the bulk band gap. Thus, one cannot distinguish the anti–localized modes and bulk modes at $b<0.3$, which implies the disappearance of the anti–localized zero modes in the linear limit.
@@
We numerically solve the nonlinear eigenvalue problem of the finite chains of the nonlinear SSH model with different system sizes $L=5$, $10$, $15$, and $20$. To solve the nonlinear eigenvalue problem, we assume the eigenequation as an algebraic equation and use the quasi–Newton method Sone2023 . Supplementary Figure S2 shows the minimum absolute values of eigenvalues and the localization indicators (Supplementary Eq. (S109)) of the corresponding eigenvectors. At $a=1.5$ (Supplementary Fig. S2a), the nonlinear winding number is nonzero $\nu_{\rm NL}=1$ in the range of amplitude, $2.50.5<w<$ 2.5. In this regime, we obtain localized zero modes corresponding to the nonzero winding number. Meanwhile, we obtain anti–localized zero modes at $w<0.5$ and no zero modes at $w>2.5$. Therefore, the nonlinear winding number predicts the existence or absence of the localized zero modes at $a=1.5$. We can also confirm the bulk–edge correspondence at $a=1.9$ as shown in Supplementary Fig. S2b.
@@
In contrast, if we consider $a=2.1$, we obtain anti–localized zero modes even in the topologically nontrivial phase at $w>1.1$ (Supplementary Fig. S2c). Therefore, the bulk–edge correspondence is broken in this parameter regime, which is consistent with the result in the semi–infinite system shown in Fig. 2 in the main text. We note that the zero modes in the semi–infinite system exhibit the localization indicator $P>1$ even at $a>2$ and thus can be regarded as locally localized but globally anti–localized zero modes. Meanwhile, the zero modes in finite systems can be locally anti–localized in the sense that the localization indicator $P$ is smaller than one. While both the breakdowns of the bulk–edge correspondence observed in finite and semi–infinite systems are related to the period–doubling bifurcation, the detailed mechanism of the anti–localization of zero modes in finite systems remains a future issue.
@@
The obtained nonlinear eigenvalues are not exactly zero due to the finite–size effect. To confirm the convergence of nonlinear eigenvalues to zero in the thermodynamic limit $L\rightarrow\infty$, we investigate the finite–size scaling of the minimum absolute values of nonlinear eigenvalues. Supplementary Figure S3a shows the minimum absolute values of nonlinear eigenvalues at different system sizes. We here use the parameters $a=1.5$, $b=c=1$, and $d=1$, and fix the edge amplitude to be $w_{\rm edge}=0.6$. We confirm that the nonlinear eigenvalues are inversely proportional to the system size and thus converge to zero in the thermodynamic limit. This system–size dependence of the nonlinear eigenvalue is different from the linear case, which is exponentially decreased in the system size. This is because the nonlinearity–induced edge modes have nonvanishing amplitudes far from the edge $x\rightarrow\pm\infty$, and the left and right localized modes $\Psi_{\rm left}$, $\Psi_{\rm right}$ have a larger resonance integral than in linear cases. In fact, the resonance integral becomes
$\frac{\sum_{x=1}^{L}\sum_{j=A,B}\Psi_{{\rm left},j}(x)^{\ast}\vec{f}(\Psi_{{\rm right},j}(x))}{\sqrt{\sum_{x=1}^{L}\sum_{j=A,B}\Psi_{{\rm left},j}(x)^{\ast}\Psi_{{\rm left},j}(x)}\sqrt{\sum_{x=1}^{L}\sum_{j=A,B}\Psi_{{\rm right},j}(x)^{\ast}\Psi_{{\rm right},j}(x)}}=\frac{\mathcal{O}(1)}{\mathcal{O}(L)}=\mathcal{O}(L^{-1}),$ (S112)
where $\vec{f}$ represents the right–hand side of Eqs. (4) and (5) in the main text. To obtain the system–size dependence, we utilize the fact that $\vec{f}(\Psi_{{\rm right},j}(x))$ is zero in the bulk, and thus
$\sum_{x=1}^{L}\sum_{j=A,B}\Psi_{{\rm left},j}(x)^{\ast}\vec{f}(\Psi_{{\rm right},j}(x))=\sum_{j=A,B}\left[\Psi_{{\rm left},j}(1)^{\ast}\vec{f}(\Psi_{{\rm right},j}(1))+\Psi_{{\rm left},j}(L)^{\ast}\vec{f}(\Psi_{{\rm right},j}(L))\right]$ (S113)
becomes almost independent of the system size. Thus, the interaction between the left and right localized modes is also proportional to the inverse of the system size, which results in the power–law system–size dependence of the nonlinear eigenvalue of edge modes. We note that similar size dependence is observed in a previous paper on the nonlinear Chern number Sone2023 .
@@
It is also noteworthy that the anti–localized zero modes disappear in the linear limit $b,c\rightarrow 0$. We consider the case of $b=c$ and numerically confirm such disappearance of the anti–localized zero modes by calculating the nonlinear eigenvalues at different $b$. Supplementary Figure S3b shows the minimum absolute values of the nonlinear eigenvalues and the size of the bulk band gaps predicted from the Bloch ansatz. At sufficiently small $b$, the nonlinear eigenvalues become larger than the bulk band gaps, and thus one cannot distinguish the anti–localized zero modes and bulk modes. Therefore, anti–localized zero modes are induced by the nonlinearity and disappear in the linear limit.
@@
Figure S4: Minimum absolute values of eigenvalues and different definitions of localization indicators $P_{1,2,3}$ of the finite system of the nonlinear SSH model. We fix the parameters $b=c=-1$ and $d=1$. We plot the minimum absolute values of the nonlinear eigenvalues at the system size $L=5$, $10$, $15$, and $20$ from the top. The red (resp. black) circles represent $P_{1,2,3}>1$ (resp. $P_{1,2,3}<1$) of the corresponding nonlinear eigenvectors. a-c Localization indicators at $a=1.5$. Panels a, b, and c show the localization indicators $P_{1}$, $P_{2}$, and $P_{3}$, respectively. The blue dashed lines are the transition point of the nonlinear winding number. We can confirm the bulk–edge correspondence between the nonlinear winding number and the localized zero modes. We note that some data points disappear at an amplitude larger than $2.5$ due to the difficulty of the convergence of the numerical techniques under the strong nonlinearity. In this parameter region, all the indicators correspond to the bulk topological invariants. d-f Localization indicators at $a=1.9$. Panels d, e, and f show the localization indicators $P_{1}$, $P_{2}$, and $P_{3}$, respectively. The blue dashed lines show the transition point of the nonlinear winding number, while the localization indicators are unchanged at those boundaries. Since the bulk–edge correspondence should still be valid in this parameter region as inferred from the bifurcation diagram in Fig. 2 in the main text, the localization indicators $P_{1,2,3}$ are useless to discuss the bulk–boundary boundary correspondence in finite systems.